OpenStudy (daniellelovee):

Propane gas (C3H8) burns completely in the presence of oxygen gas (O2) to yield carbon dioxide gas (CO2) and water vapor (H2O). • Assuming that all volume measurements occur at the same temperature and pressure, how many liters of carbon dioxide gas will be produced in the reaction?

OpenStudy (daniellelovee):

@Somy @ichliebedich @SyedMohammed98 the crew lol

OpenStudy (ichliebedich):

C3H8 + 5 O2 → 3 CO2 + 4 H2O = (0.700 L C3H8) x (3 mol CO2 / 1 mol C3H8) (0.700 L C3H8) x (4 mol H2O / 1 mol C3H8) = 2.80 L H2O vapor

OpenStudy (daniellelovee):

i've got that already but that is the answer to the previous questions...

OpenStudy (ichliebedich):

this one? (0.700 L C3H8) X (5 mol O2 / 1 mol C3H8) = 3.50 L O2

OpenStudy (daniellelovee):

no that is the answer for the previous question it has 3 measurements but aIready got that I need this question :/

OpenStudy (somy):

is it done? + 0.700 is a given value?

OpenStudy (daniellelovee):

would this be the answer then? (0.700 L C3H8) x (3 mol CO2 /1mol C3H8) = 2.10 L CO2

OpenStudy (somy):

so 0.700 is given yeah it is correct

OpenStudy (somy):

its simple math game \(C_3H_8 + 5O_2 → 3CO_2 + 4H_2O\) 1 - 0.700 3 - X X= 2.10 L Of CO2

OpenStudy (daniellelovee):

alright thank you for like the 5th time today xD

OpenStudy (somy):

:))

OpenStudy (ichliebedich):

sorry im not that much of a help

OpenStudy (daniellelovee):

nah you are don't worry ^.^

OpenStudy (somy):

@ichliebedich you are more of a help than me lol

OpenStudy (ichliebedich):

thanks @Daniellelovee and @Somy :)

OpenStudy (daniellelovee):

no problem :)

OpenStudy (somy):

hehe :)